# Analog charged black hole formation via percolation: Exploring cosmic censorship and Hoop conjecture

### v1

@@ I Introduction
Black holes (BHs) pose a challenge to our understanding of the fundamental laws of physics [1, 2, 3]. The BH information paradox, stemming from the apparent conflict between quantum mechanical unitarity and the seemingly irreversible nature of BH evaporation [2, 3], highlights a critical gap in our knowledge. This paradox hinges critically on our understanding of the nature of singularities [4] and their fate in quantum gravity [5, 6, 7, 8, 9, 10, 11, 12]. The cosmic censorship hypothesis, if true, suggests that these singularities are always cloaked by event horizons, further complicating the question of how information might escape [13]. While a complete theory of quantum gravity is expected to provide the ultimate resolution [14], its development remains a formidable challenge. A central question is whether the paradox can be resolved without a full understanding of singularities [15, 16, 17, 18].
Condensed matter systems provide a fertile ground for exploring this enigma. These systems exhibit phenomena remarkably analogous to BHs, including quasiparticle emergence, topological phase transitions, and complex entanglement dynamics [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. For instance, the emergence of quasiparticles from collective behavior mirrors Hawking radiation from a BH horizon [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. By studying these systems, we can explore how information is processed in strongly correlated systems [42, 43, 44, 45], gaining insights into the mechanisms of information preservation and loss [46, 47, 48, 49, 50, 51]. While these systems are BH analogs, not exact replicas, they provide valuable experimental platforms for testing theoretical ideas and deepening our understanding of BHs [52, 53, 54, 55, 56, 57].
@@
Chirality plays a crucial role in condensed matter physics, as exemplified by the chiral magnetic effect in Weyl semimetals[59, 60, 61]. In our model, the chiral term induces an imbalance in spin excitations, analogous to the charge imbalance in the chiral magnetic effect, which, in conjunction with the inhomogeneities of the non-chiral term, drives the system toward a topological QPT [58]. Chiral spin-chain models are also relevant to understanding non-perturbative phenomena in quantum field theory via lattice techniques [62, 63]. However, lattice formulations of chiral theories face the challenge of fermion doubling, as dictated by the Nielsen-Ninomiya theorem [64].
By incorporating NNN interactions, our model effectively circumvents the doubling problem. The NNN interacting chiral term ($\chi_{j}=\vec{\sigma}_{j}\cdot\left(\vec{\sigma}_{j+1}\times\vec{\sigma}_{j+2}\right)$) acts as a long-range interaction, despite being constructed from local terms. This arises because the chirality operator effectively couples spins across multiple lattice sites, creating long-range correlations. The expectation value of the chirality operator, $\braket{\chi_{j}}$, encodes information about the system’s global topology [65, 66, 67]. These long-range correlations bypass Nielsen-Ninomiya theorem [64].
The formation of BHs has been linked to critical phenomena, exhibiting power-law scaling near the threshold of BH formation, analogous to phase transitions, where fluctuations occur at all length scales due to the divergence of the correlation length [68]. This work offers a novel perspective, connecting BH formation to percolation theory, a powerful framework for understanding cluster formation and connectivity at criticality, where large-scale connectivity emerges, modeling diverse physical phenomena [69]. The key idea is when matter collapses under gravity, fluctuations occur at different length scales leading to regions of varying energy density. When both the density of the collapsing region and the number of connected clusters surpass a critical threshold, these regions percolate, leading to the formation of an event horizon. Specifically, we demonstrate a correspondence between the percolation threshold in a spin chain model [69, 70] and the emergence of an event horizon-like boundary, defined by NNN-dominated regions (interior) versus nearest-neighbor (NN)-dominated regions (exterior). Within the NNN-dominated regime, we observe exponential growth in both interconnected spin clusters (analogous to matter aggregation) and the energy density within these clusters (akin to BH formation). Furthermore, we find that the critical exponents associated with the total cluster energy and the percolation correlation length correspond to the BH mass and surface gravity exponents, respectively. We discuss the implications for cosmic censorship and the hoop conjecture.
@@ III Analog BH
The model admits a geometric interpretation in the context of BH physics. To establish this connection, we have assigned the lattice sites of (5) alternately to sublattices $A$ and $B$, which facilitates the emergence of Dirac-like quasiparticles in the continuum limit with the following 2-D metric (see Appendix B)
@@
where $d\mathcal{X}={dx}/{\mathcal{U}(x)}$ and $\mathcal{U}(x)$ is arbitrary. To further analyze, we take a specific inhomogeneous form for $\mathcal{U}(x)$:
@@
where $\sigma$ is a constant. Thus, for $\sigma<\sigma_{c}\equiv{1}/{(2v^{2})}$, the $\tau-\tau$ component of the metric (10) vanishes at two distinct points. This condition signifies the presence of two horizons, defined by $g(x)=0$ and located at:
@@
To further understand the connection between the analog model and RN BH, we evaluate three geometric quantities for the analog metric (8): the Kretschmann scalar ($\mathcal{K}$), surface gravity ($\kappa_{g}$), and Komar mass $(M_{\rm BH}$). $\mathcal{K}$ provides information about the curvature of the space-time and is $\mathcal{K}=1/(\sigma^{4}\sigma_{c}^{2})$. Depending on $v$, $0<\mathcal{K}<\infty$. The surface gravity and Komar mass are
$\kappa_{g}^{2}$ $=$ $\frac{1}{8\sigma\sigma_{c}^{2}}\left[6\sinh^{2}\left(\frac{x}{\sigma}\right)-\frac{1}{\sigma}\tanh^{2}\left(\frac{x}{\sigma}\right)g^{2}(x)\right]~{},$ (14)
$M_{BH}$ $=$ $\frac{1}{\sigma_{c}}\sqrt{\frac{2}{\sigma}}\tan^{-1}\left[\tanh\left(\frac{1}{2}\cosh^{-1}\left[\sqrt{\frac{\sigma_{c}}{{\sigma}}}\right]\right)\right].~{}~{}~{}~{}~{}$ (15)
Later, these quantities will be employed to investigate the nature of the phase transition in the spin-chain and determine whether it exhibits characteristics analogous to a RN BH. In Appendix B, we discuss limits when $v$ or ${\cal U}$ vanish or diverge, leading to Carrollian and shockwave limits.
@@ IV.1 Percolation exponents
where IE is $\mathcal{U}$ for NN and $v$ for NNN interactions. $\rho_{T}$ is obtained by summing the energy densities of all individual clusters. As shown in Fig. 2, we see two distinct features: When NN interactions dominate, the normalized total energy density of the clusters $\rho_{T}/\mathcal{U}$, shown in red (left) and ${\cal N}$ shown in red (right) increases linearly with $N$, indicating a simple connectivity structure. However, when NNN interactions dominate, the normalized total energy density of the clusters $\rho_{T}/v$ shown in blue (left) and the number of clusters shown in blue (right) exhibit exponential growth, characteristic of a BH-like phase with intricate connectivity. The contrasting growth provides strong evidence for BH phase interpretation.
@@ V Conclusions
This percolation framework describes gravitational collapse and BH formation. The emergence of an event horizon is analogous to reaching the percolation threshold, with distinct regions inside and outside the horizon reflecting the transition between physical phases. The close alignment of scaling exponents derived from the analog model with those of surface gravity and Komar mass in the BH context further validate this framework. Crucially, the exponential cluster growth observed in analogous to matter aggregation during stellar collapse.
@@ Acknowledgments
The authors thank G. Baskaran, Indranil Chakraborty, K. Hari, P. George Christopher, Susobhan Mandal, Tausif Parvez, Tapobrata Sarkar, and Semin Xavier for discussions and comments on the earlier draft. The work is supported by SERB-CRG (RD/0122-SERB000-044).
@@ Appendix B The continuum limit and emergent black hole background
$\tilde{E}(p)=\pm\,\mathcal{V}\left(\mathcal{U}\sqrt{2+2\cos(a_{c}p)}+v\sin(a_{c}p)\right).$

### v2

@@ I Introduction
Black holes (BHs) pose challenges to our understanding of fundamental physics [1, 2, 3]. The BH information paradox, stemming from the apparent conflict between quantum mechanical unitarity and the seemingly irreversible nature of BH evaporation [2, 3], highlights a critical gap in our knowledge, intertwined with our understanding of the nature of singularities [4] and their fate in quantum gravity [5, 6, 7, 8, 9, 10, 11, 12]. The cosmic censorship hypothesis further complicates this by suggesting singularities are always hidden by event horizons [13]. While quantum gravity is anticipated to provide a resolution [14], its development remains a major hurdle, prompting investigations into whether the paradox can be addressed without a complete theory of singularities [15, 16, 17, 18].
Condensed matter systems provide a fertile ground for exploring this enigma. These systems exhibit phenomena remarkably analogous to BHs, including quasiparticle emergence, topological phase transitions, and complex entanglement dynamics [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. The emergence of quasiparticles, for instance, mirrors Hawking radiation [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. Studying information processing in these strongly correlated systems [42, 43, 44, 45] can provide insights into information preservation and loss [46, 47, 48, 49, 50, 51]. While these are analogs, not exact replicas, they serve as crucial experimental platforms for testing theoretical ideas [52, 53, 54, 55, 56, 57, 58, 59].
@@
Chirality plays a crucial role in condensed matter physics, as exemplified by the chiral magnetic effect in Weyl semimetals [61, 62, 63]. In our model, the chiral term induces an imbalance in spin excitations, analogous to the charge imbalance in the chiral magnetic effect, which, in conjunction with the inhomogeneities of the non-chiral term, drives the system toward a topological QPT [60]. Chiral spin chains are also relevant for non-perturbative QFT via lattice techniques [64, 65]. Lattice chiral theories face fermion doubling issues [66], which our model circumvents using NNN interactions ($\chi_{j}=\vec{\sigma}_{j}\cdot\left(\vec{\sigma}_{j+1}\times\vec{\sigma}_{j+2}\right)$). This term acts as a long-range interaction, despite being constructed from local terms. This arises because the chirality operator effectively couples spins across multiple lattice sites, creating long-range correlations. The expectation value of the chirality operator, $\braket{\chi_{j}}$, encodes information about the system’s global topology [67, 68, 69]. These long-range correlations bypass Nielsen-Ninomiya theorem [66].
The formation of BHs has been linked to critical phenomena, exhibiting power-law scaling near the threshold of BH formation, analogous to phase transitions with diverging correlation length [70]. A key geometric criterion for predicting horizon formation in diverse scenarios, including non-symmetric collapse, is the hoop conjecture [71], which postulates that a mass ${\cal M}$ collapses to form a BH if and only if it can be enclosed by a circular hoop with a critical circumference ${\cal C}_{\rm critical}=4\pi{\cal M}$ [72, 73, 74], implying ${\cal C}\leq 4\pi{\cal M}$. This suggests that sufficient mass concentration is crucial for horizon formation.
This work offers a novel perspective, connecting BH formation to percolation theory, a powerful framework for understanding cluster formation and connectivity at criticality, where large-scale connectivity emerges, modeling diverse physical phenomena [75]. We propose that gravitational collapse involves fluctuations leading to varying energy densities, and when both the density and connected clusters surpass a critical threshold, percolation occurs, forming an event horizon. Specifically, we demonstrate a correspondence between the percolation threshold in a spin chain model [75, 76] and the emergence of an event horizon-like boundary, defined by NNN-dominated regions (interior) versus nearest-neighbor (NN)-dominated regions (exterior). Within the NNN-dominated regime, we observe exponential growth in both interconnected spin clusters (analogous to matter aggregation) and the energy density within these clusters (akin to BH formation). Furthermore, we find that the critical exponents associated with the total cluster energy and the percolation correlation length correspond to the BH mass and surface gravity exponents, respectively. We discuss the implications for cosmic censorship and the hoop conjecture.
@@ III Analog BH
The model admits a geometric interpretation in the context of BH physics. To establish this connection, we partitioned the lattice sites of $H_{MF}$ in (5) into alternating sublattices A and B, a procedure that facilitates the emergence of Dirac-like quasiparticles upon taking the continuum limit. This limit is achieved by expanding the mean-field Hamiltonian using a Taylor series around the Fermi points, where the characteristic Dirac cone structure arises. In the low-energy regime, the linear component of the dispersion relation dominates, while higher-order terms are negligible. This approximation yields an effective Dirac Hamiltonian defined on a 2-D spacetime with the following metric (details in Appendix B):
@@
where $d\mathcal{X}={dx}/{\mathcal{U}(x)}$. Until this point, ${\mathcal{U}(x)}$ is arbitrary, however, to establish a meaningful correspondence with an analog BH solution, we need to make a specific choice of ${\mathcal{U}(x)}$. In the rest of this work, we choose the form of $\mathcal{U}(x)$ that is smooth, and asymptotically decays:
@@
where $\sigma$ is a constant. For $\sigma<\sigma_{c}\equiv{1}/{(2v^{2})}$, the $\tau-\tau$ component of the metric (10) vanishes at two distinct points. This condition signifies the presence of two horizons, defined by $g(x)=0$ and located at:
@@
To further understand the connection between the analog model and RN BH, we evaluate three geometric quantities for the analog metric (8): the Kretschmann scalar ($\mathcal{K}$), surface gravity ($\kappa_{g}$), and Komar mass $(M_{K}$[87]. $\mathcal{K}$ is a quadratic scalar invariant of the Riemann tensor [87, 88], defined as $\mathcal{K}=R_{\mu\nu\rho\eta}R^{\mu\nu\rho\eta}$. For metric (8), $\mathcal{K}=1/(\sigma^{4}\sigma_{c}^{2})$. Depending on $v$, $0<\mathcal{K}<\infty$. The line-element (8) has a timelike Killing vector $K^{\mu}$, $\kappa_{g}$ at event horizon is defined as $\kappa_{g}^{2}=-\frac{1}{2}(\nabla_{\mu}K_{\nu})(\nabla^{\mu}K^{\nu})$:
The Komar mass is [88],
[…]
where $S$ is a constant time slice, $g$ is the determinant of the metric, and $n^{\mu}$ is a unit vector orthogonal to $S$. The boundary of $S$ has two contributions, one from an inner boundary of the horizon which is BH mass evaluated as
and the second from an outer boundary at infinity. Later, these quantities will be employed to investigate the nature of the phase transition in the spin-chain and determine whether it exhibits characteristics analogous to a RN BH. In Appendix B, we discuss limits when $v$ or ${\cal U}$ vanish or diverge, leading to Carrollian and shockwave limits.
@@ IV.1 Percolation exponents
where “# of connected states” refers to the distinct states within a given cluster that are connected via one or more bonds and IE is $\mathcal{U}$ for NN and $v$ for NNN interactions. $\rho_{T}$ is obtained by summing the energy densities of all individual clusters. As shown in Fig. 2, we see two distinct features: When NN interactions dominate, the normalized total energy density of the clusters $\rho_{T}/\mathcal{U}$, shown in red (left) and ${\cal N}$ shown in red (right) increases linearly with $N$, indicating a simple connectivity structure. However, when NNN interactions dominate, the normalized total energy density of the clusters $\rho_{T}/v$ shown in blue (left) and the number of clusters shown in blue (right) exhibit exponential growth, characteristic of a BH-like phase with intricate connectivity. The contrasting growth provides strong evidence for BH phase interpretation.
@@ V Conclusions
Classical percolation effectively models how local connectivity rules generate macroscopic structures and critical behavior near transitions. This directly parallels black hole formation, accurately reproducing critical exponents. The percolating cluster emergence mirrors event horizon formation, where increasing interaction strength analogously represents growing gravitational influence. The close alignment of scaling exponents derived from the analog model with those of surface gravity and Komar mass in the BH context further validate this framework. Crucially, the exponential cluster growth observed in analogous to matter aggregation during stellar collapse.
@@ Acknowledgments
The authors thank G. Baskaran, Indranil Chakraborty, K. Hari, P. George Christopher, Susobhan Mandal, Tausif Parvez, Tapobrata Sarkar, and Semin Xavier for discussions and comments on the earlier draft. We also thank the authors of Ref. [68] for email correspondence. The work is supported by SERB-CRG (RD/0122-SERB000-044).
@@ Appendix B The continuum limit and emergent black hole background
$\tilde{E}(p)=\mathcal{V}\left(\pm\,\mathcal{U}\sqrt{2+2\cos(a_{c}p)}+v\sin(a_{c}p)\right).$